OpenStudy (anonymous):

The area of Greenland is 840,000 square miles. What is this number in scientific notation? I have no idea what scientific notation is!!

OpenStudy (anonymous):

@Hero

hero (hero):

You convert it to \(8.4 \times 10^{5}\)

OpenStudy (anonymous):

where did you get \[10^{5}\]

hero (hero):

Every number is associated with a decimal point. We don't write decimal points with every number because it is convenient not to. 840,000 can easily be re-written as 840000.0

hero (hero):

If you move the decimal point 5 places to the left, you'll have 8.40000

OpenStudy (anonymous):

ok so you got \[10^{5}\] by the zeros

hero (hero):

But \(8.40000 = 8.4 \times 100000\) And \(100000 = 10^5\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what is scientific notation?

hero (hero):

\(8.4 \times 10^5\) is scientific notation. In other words a decimal number multplied by base 10 times an exponent.

OpenStudy (anonymous):

ohhhh ok that makes sense